# Scrapers selectively respect robots.txt directives: evidence from a large-scale empirical study

### v1

@@ Abstract.
Online data scraping has taken on new dimensions in recent years, as traditional scrapers have been joined by new AI-specific bots. To counteract unwanted scraping, many sites use tools like the Robots Exclusion Protocol (REP), which places a robots.txt file at the site root to dictate scraper behavior. Yet, the efficacy of the REP is not well-understood. Anecdotal evidence suggests some bots comply poorly with it, but no rigorous study exists to support (or refute) this claim. To understand the merits and limits of the REP, we conduct the first large-scale study of web scraper compliance with robots.txt directives using anonymized web logs from our institution. We analyze the behavior of $130$ self-declared bots (and many anonymous ones) over $40$ days, using a series of controlled robots.txt experiments. We find that bots are less likely to comply with stricter robots.txt directives, and that certain categories of bots, including AI search crawlers, rarely check robots.txt at all. These findings suggest that relying on robots.txt files to prevent unwanted scraping is risky and highlight the need for alternative approaches.
@@ 1. Introduction
Beyond well-documented copyright and privacy concerns, widespread scraping for any purpose—including for AI training data, RAG, and agent use—can destabilize public, data-rich websites. Numerous writings from the “grey” (non-academic) literature describe sites being taken offline in recent months due to thousands of download requests from scrapers associated with AI companies (Mehrotra and Marchman, 2024; Iaso, 2025; Claburn, 2024). These scraping-induced problems have prompted an escalating tactical war, as content-rich websites attempt to moderate or prevent unwanted web scraping (Belanger, 2025; Longpre et al., 2024; Longpre, 2025).
Among the most widely used anti-scraping solution is the Robots Exclusion Protocol (REP), implemented via a robots.txt file located at the site root. The REP lets site owners set rules that specify which bots can scrape information from the site, which sub-domains those bots can access, and how long bots must wait between successive page accesses (rep, 2022). Recent work (Longpre et al., 2024) showed a significant uptick in the use of and restrictions in robots.txt files after the rise of generative AI models around 2022, purportedly in reaction to increased scraping for training data. The REP was codified in 2022, but it is not legally binding, so compliance requires scraper goodwill.
@@ 2.2. Bot deterrence
Although certain types of web scraping can benefit site owners (e.g. search engine optimization), other types can be harmful. For example, unfettered scraping can scoop up data that, while publicly available, may have copyright or privacy concerns (Longpre et al., 2024). Furthermore, numerous blog posts and other articles from recent years have documented how high-intensity scraping by presumed AI crawlers have caused site instability (Mehrotra and Marchman, 2024; Iaso, 2025; Belanger, 2025). In light of these downsides, many techniques have been proposed to prevent unwanted scraping.
@@
User-agent: Googlbot
@@
The Robots Exclusion Protocol emerged in 1994 as a possible solution to unwanted bot activity. As crawlers became prominent on the early web, it became clear that a mechanism was needed to allow “friendly” scrapers and disallow poorly-behaved ones (Koster, 1994). The Robots Exclusion Protocol (REP) (rep, 2022) was proposed to fill this gap. The REP asks developers to place a file called robots.txt at their site root, specifying restrictions on certain bots and bot behaviors on their site. Possible bot directives include restricting subdomain access, enforcing a crawl delay, and allowing only bots with certain user-agent strings—see Table 1 and Figure 1. The REP was codified as RFC 9309 by the Internet Engineering Task Force (IETF) (Force, 2022) and is widely used (Sun et al., 2007; Longpre et al., 2024).
Despite its prominence, the REP has significant drawbacks. First, robots.txt is not a legally binding document, so requires the good-will of the parties involved for compliance (rbt, 2007). Second, it requires web hosts to maintain extensive knowledge of user agents and entities that they wish to exclude from scraping (Longpre et al., 2024). This places a heavy burden on hosts to maintain awareness of the ever-changing web landscape. Third, since compliance is optional, some crawlers simply ignore robots.txt . Finally, limited work (Liu et al., 2024) has studied bot compliance with robots.txt , and none have done so at scale, making its efficacy unknown.
Other bot-blocking methods. Other methods to block unwanted scraping behavior vary from mildly intrusive to very disruptive. CAPTCHAs, which ask site visitors to solve a hard (e.g. computationally difficult) problem before viewing or retrieving page content, have historically been effective in deterring bots (Von Ahn et al., 2003, 2008). However, they add significant friction to user interactions with webpages (Bursztein et al., 2010), while recent advances in AI make CAPTCHA solving easy (Bursztein et al., 2014; Guerar et al., 2021; Ye et al., 2018), both of which make CAPTCHAs less useful. Companies like Cloudflare market proprietary bot detection and deterrence methods, which can be purchased (Cloudflare, 2024). As a last resort, companies can outright block the IP addresses of troublesome bots. However, since many bots run on VPNs, IP addresses can be easily recycled, circumventing this approach. A recent work suggests using reverse proxies to better regulate traffic to a site (Liu et al., 2024). Finally, other articles suggest novel methods like employing a proof of work (Iaso, 2025) or even a tarpit (Belanger, 2025), which creates unending fake content for scrapers.
@@ 3.1. Dataset Preparation
To study scraper compliance with robots.txt files at scale, we collect and analyze a large dataset composed of network access data from website endpoints managed by our institution. The dataset contains approximately 3.9 million web requests made to a set of 36 websites from February 12 - March 29, 2025. The data include the following fields:
  • Useragent: A string identifying the entity accessing a website that often includes the browser, operating system, hardware, and version information of the origin of the request. Well-known bots often provide a simpler identifying string in this field (e.g. Googlebot).
@@
  • URI path: The subdomain accessed within the website. The sitename and URI path combined form the whole URL.
@@
  • Bytes: Bytes collected by web visitor during site interaction (e.g. amount of downloaded content).
  • Referer: Site from which web visitor was redirected.
@@
Table 3. Over $40\%$ of web accesses in our dataset are attributable to just $20$ bots. We document the number of unique web accesses for each bot, the proportion of web traffic they compose, and the amount of data each scraped during the 40 day period.
@@
Figure 3. Most bot in the top $5$ categories in terms of data scraped collect data steadily, but search engine crawlers buck the trend, driven by YisouSpider’s mid-March activity. AI assistants scrape much data relative to session count (Fig 4).
@@
Data preprocessing. Data from any logged-in institutional users is filtered out before it reaches our servers to protect user privacy. We inspect data after collection and remove several IP hashes associated with internal vulnerability scanning tools and similar entities that are not relevant to our analysis. We screen out $3$ IP hashes associated with $294,362$ web accesses through this approach. Then, we map ASNs to ARIN info and standardize bot names to enhance the data and simplify analysis.
To map ASNs to ARIN info, we leverage the external library whoisit [^1] https://pypi.org/project/python-whois/ to poll for whois information for all unique ASNs in our incoming dataset. The returned whois information contains information held by ARIN about the registered user behind the ASN, including the declared entity name and description. We also standardize bot names via fuzzy string matching with a public dataset of common useragent strings and the corresponding bot names [^2] https://github.com/monperrus/crawler-user-agents. Using these standardized names, we add an additional column to our dataset that maps bots to bot categories proposed by Dark Visitors, an industry website that keeps an up-to-date list of known scrapers [^3] https://darkvisitors.com/—AI Agents, AI Assistants, AI Data Scrapers, Archivers, Developer Helpers, Fetchers, Headless Agents, Intelligence Gatherers, Scrapers, Search Engine Crawlers, SEO Crawlers, Uncategorized, and Undocumented AI Agents. This categorization allows us to analyze collective bot behaviors.
Ethical considerations. Data collection and analysis adhered to a study protocol approved by our local Institutional Review Board (IRB) [^4] IRB 2025-0094. We followed all IRB guidelines for anonymizing and storing this dataset, including making significant efforts to exclude non-bot behavior from the dataset and preserve user privacy. We removed sessions originating from logged-in users from our institution or from institution-hosted WiFi networks, and anonymized IP addresses by hashing them. All data was stored and analyzed on secure servers only accessible by the research team.
@@ 3.2. Dataset overview
Most active bots. Now, we focus specifically on known bots, e.g. bots with self-declared user agents that are well-documented by several web sources. Table 3 gives details about the behavior of the 20 most active bots (as measured by number of web accesses) seen in our dataset. Together, these $20$ bots make up over $40\%$ of the observed web traffic. YisouSpider and AppleBot dominate, collectively driving $30\%$ of all traffic and scraping over $8$ GB of data during the 40 day period. YisouSpider is associated with the Chinese search engine Yisou and appears to be used to index web pages (Udger, 2025). AppleBot is a generic bot associated with Apple and is “used by Apple to index search results that allow the Siri AI Assistant to answer user questions” (Apple, 2025).
@@ 4.1. Methodology
Over a period of 8 weeks, we deployed four versions of robots.txt files on one site controlled by our institution. This site was chosen because of its observed high bot traffic, making it a rich source of data on scraper behavior. Each robots.txt file was deployed for two weeks, then swapped out by support staff for the next version. The four versions of the robots.txt files have increasingly strict directives, ranging from no meaningful restrictions to full denial of access for certain bots. Using multiple versions of robots.txt allowed us to achieve two distinct goals: measuring whether certain robots.txt directives were more or less respected by scrapers, and measuring how quickly scrapers adapted to new robots.txt restrictions. The files are described below and shown in Figures 8-8.
@@
  • Version 1 (Fig. 8): All bots retain the same level of access as in the base version, but a 30 second crawl delay is requested between successive page requests. Crawl delay directive is not part of the official robots.txt protcol (rep, 2022) but is widely used, making it a useful directive to study.
  • Version 2 (Fig. 8): Most bots are allowed to access only the /page-data/* endpoint, which we observed experimentally to be a common target for scrapers, but are disallowed from all other sites. Eight SEO bots [^5] Googlebot, Slurp, bingbot, Yandexbot, DuckDuckBot, BaiduSpider, DuckAssistBot, ia_archiver are exempt from these restrictions per our institution’s request.
@@
Data preparation. We performed basic preprocessing on our datasets after collection to aid analysis. Much of our preprocessing pipeline for the robots.txt experiments mirrored that described in §3. However, for our analysis of compliance with robots.txt we also filter out bots that accessed the site less than $5$ times under any robots.txt version.
@@
AI Assistants 0.910 (2077) 0.131 (2154) 1.000 (1126) 0.665
AI Data Scrapers 0.560 (1862) 0.352 (1824) 0.771 (1822) 0.560
AI Search Crawlers 0.895 (1157) 0.673 (975) 0.348 (1056) 0.644
Fetchers 0.927 (252) 0.277 (252) 0.377 (235) 0.525
Headless Browsers 0.036 (2777) 0.278 (4207) 0.011 (1282) 0.153
Intelligence Gatherers 0.809 (1120) 0.372 (549) 0.094 (1008) 0.431
Other 0.486 (5687) 0.119 (4506) 0.016 (4770) 0.220
SEO Crawlers 0.645 (1580) 0.832 (1217) 0.639 (1268) 0.704
Search Engine Crawlers 0.777 (6379) 0.365 (5934) 0.186 (6409) 0.443
Directive average 0.618 0.314 0.261 0.397
@@
Endpoint access compliance (v2). This robots.txt version restricts bots (except for a group of SEO bots, see §4.1) to only accessing subpages under a single endpoint, /page-data, on the site. Computing the ratio of compliance with this directive is more straightforward. For each unique user-agent, we count its accesses to either the robots.txt file (which is always allowed) or /page-data subpages and compute the ratio of this to the user agent’s total page accesses. This compliance ratio $C_{\tau_{i}}$ is $1$ if the user-agent only accesses the allowed pages, robots.txt and page-data, and is near 00 if it consistently accesses many other pages. We then compute an identical compliance metric for user agents under our default robots.txt and use this as the baseline for a paired z-test for differences in proportions between the two datasets. Ideally, we should see a significant uptick in compliance when the endpoint directive is in place.

### v2

@@ Abstract.
Online data scraping has taken on new dimensions in recent years, as traditional scrapers have been joined by new AI-specific bots. To counteract unwanted scraping, many sites use tools like the Robots Exclusion Protocol (REP), which places a robots.txt file at the site root to dictate scraper behavior. Yet, the efficacy of the REP is not well-understood. Anecdotal evidence suggests some bots comply poorly with it, but no rigorous study exists to support (or refute) this claim. To understand the merits and limits of the REP, we conduct the first large-scale study of web scraper compliance with robots.txt directives using anonymized web logs from our institution. We analyze the behavior of $130$ self-declared bots (and many anonymous ones) over $40$ days, using a series of controlled robots.txt experiments. We find that bots are less likely to comply with stricter robots.txt directives, and that certain categories of bots, including AI search crawlers, rarely check robots.txt at all. Our findings suggest that relying on robots.txt to prevent unwanted scraping is risky and highlight the need for alternatives.
web scraping; content controls; bots
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM Internet Measurement Conference; October 28–31, 2025; Madison, WI, USAbooktitle: Proceedings of the 2025 ACM Internet Measurement Conference (IMC ’25), October 28–31, 2025, Madison, WI, USAdoi: 10.1145/3730567.3764471isbn: 979-8-4007-1860-1/2025/10ccs: Information systems Web log analysisccs: Security and privacy Access control
@@ 1. Introduction
Beyond well-documented copyright and privacy concerns, widespread scraping for any purpose—including for AI training data, RAG, and agent use—can destabilize public, data-rich websites. Numerous writings from the “grey” (non-academic) literature describe sites being taken offline in recent months due to thousands of download requests from scrapers associated with AI companies (Mehrotra and Marchman, 2024; Iaso, 2025; Claburn, 2024). These scraping-induced problems have prompted an escalating tactical war, as content-rich websites attempt to moderate or prevent unwanted web scraping (Belanger, 2025; Longpre et al., 2025; Longpre, 2025).
Among the most widely used anti-scraping solution is the Robots Exclusion Protocol (REP), implemented via a robots.txt file located at the site root. The REP lets site owners set rules that specify which bots can scrape information from the site, which sub-domains those bots can access, and how long bots must wait between successive page accesses (rep, 2022). Recent work (Longpre et al., 2025) showed a significant uptick in the use of and restrictions in robots.txt files after the rise of generative AI models around 2022, purportedly in reaction to increased scraping for training data. The REP was codified in 2022, but it is not legally binding, so compliance requires scraper goodwill.
@@ 2.2. Bot deterrence
Although certain types of web scraping can benefit site owners (e.g. search engine optimization), other types can be harmful. For example, unfettered scraping can scoop up data that, while publicly available, may have copyright or privacy concerns (Longpre et al., 2025). Furthermore, numerous blog posts and other articles from recent years have documented how high-intensity scraping by presumed AI crawlers have caused site instability (Mehrotra and Marchman, 2024; Iaso, 2025; Belanger, 2025). In light of these downsides, many techniques have been proposed to prevent unwanted scraping.
@@
User-agent: Googlebot
@@
The Robots Exclusion Protocol emerged in 1994 as a possible solution to unwanted bot activity. As crawlers became prominent on the early web, it became clear that a mechanism was needed to allow “friendly” scrapers and disallow poorly-behaved ones (Koster, 1994). The Robots Exclusion Protocol (REP) (rep, 2022) was proposed to fill this gap. The REP asks developers to place a file called robots.txt at their site root, specifying restrictions on certain bots and bot behaviors on their site. Possible bot directives include restricting subdomain access, enforcing a crawl delay, and allowing only bots with certain user-agent strings—see Table 1 and Figure 1. The REP was codified as RFC 9309 by the Internet Engineering Task Force (IETF) (Force, 2022) and is widely used (Sun et al., 2007; Longpre et al., 2025).
Despite its prominence, the REP has significant drawbacks. First, robots.txt is not a legally binding document, so requires the good-will of the parties involved for compliance (rbt, 2007). Second, it requires web hosts to maintain extensive knowledge of user agents and entities that they wish to exclude from scraping (Longpre et al., 2025). This places a heavy burden on hosts to maintain awareness of the ever-changing web landscape. Third, since compliance is optional, some crawlers simply ignore robots.txt . Finally, limited work (Liu et al., 2024) has studied bot compliance with robots.txt , and none have done so at scale, making its efficacy unknown.
Other bot-blocking methods. Other methods to block unwanted scraping behavior vary from mildly intrusive to very disruptive. CAPTCHAs, which ask site visitors to solve a hard (e.g. computationally difficult) problem before viewing or retrieving page content, have historically been effective in deterring bots (Von Ahn et al., 2003, 2008). However, they add significant friction to user interactions with webpages (Bursztein et al., 2010), while recent advances in AI (and human skill) make CAPTCHA solving easy (Bursztein et al., 2014; Guerar et al., 2021; Ye et al., 2018; Amin Azad et al., 2020), both of which make CAPTCHAs less useful. Companies like Cloudflare market proprietary bot detection and deterrence methods, which can be purchased (Cloudflare, 2024). As a last resort, companies can outright block the IP addresses of troublesome bots. However, since many bots run on VPNs, IP addresses can be easily recycled, circumventing this approach. Several academic works suggest using reverse proxies to better regulate traffic to a site (Liu et al., 2024; Dai Nguyen et al., 2024). Finally, other articles suggest novel methods like employing a proof of work (Iaso, 2025) or even a tarpit (Belanger, 2025), which creates unending fake content for scrapers.
@@ 3.1. Dataset Preparation
To study scraper compliance with robots.txt files at scale, we collect and analyze a large dataset composed of network access data from website endpoints managed by our institution, a large private US university. The sites host information about a variety of university functions, from the IT department to campus dining to a personnel directory and beyond. The dataset contains approximately 3.9 million external web requests made to a set of 36 websites from February 12 - March 29, 2025 and includes the following fields:
  • Useragent: A self-reported string identifying the entity accessing a website that often includes the browser, operating system, hardware, and version information of the origin of the request. Well-known bots often provide a simpler identifying string in this field (e.g. Googlebot).
@@
  • URI path: The requested resource. The sitename and URI path combined form the whole URL.
@@
  • Bytes: Amount of data, in bytes, transmitted by the web server during site interaction (e.g. amount of downloaded content).
  • Referer: Site from which web visitor was redirected (if applicable.)
@@
Table 3. Over $40\%$ of web accesses in our dataset are attributable to just 2020 bots. We document the number of unique web accesses for each bot, the proportion of web traffic they compose, and the amount of data each scraped during the 40 day period.
@@
Figure 3. Most bots in the top $5$ categories in terms of data scraped collect data steadily, but search engine crawlers buck the trend, driven by YisouSpider’s mid-March activity. AI assistants scrape much data relative to session count (Fig 4).
@@
Data preprocessing. Data from any institution-hosted networks or logged-in institutional users is filtered out before it reaches our servers to protect user privacy. We inspect data after collection and remove several IP hashes associated with vulnerability scanning tools and similar entities that are not relevant to our analysis. We screen out $3$ IP hashes associated with $294,362$ web accesses through this approach. Then, we map ASNs to ARIN info and standardize bot names to enhance the data and simplify analysis.
To map ASNs to ARIN info, we leverage the external library whoisit [^1] https://pypi.org/project/python-whois/ to poll for whois information for all unique ASNs in our incoming dataset. The returned whois information contains information held by ARIN about the registered user behind the ASN, including the declared entity name and description. We also standardize bot names via fuzzy string matching with a public dataset of common useragent strings and the corresponding bot names [^2] https://github.com/monperrus/crawler-user-agents. Using these standardized names, we add an additional column to our dataset that maps bots to bot categories proposed by Dark Visitors, an industry website that keeps an up-to-date list of known scrapers [^3] https://darkvisitors.com/—AI Agents (bots from AI companies with “agent” in their name, presumed to operate as part of an agent pipeline), AI Assistants (bots that retrieve content to supplement AI queries), AI Data Scrapers (bots that scrape AI training data), Archivers, Developer Helpers, Fetchers, Headless Agents, Intelligence Gatherers (data collection for non-SEO or AI purposes), Scrapers, Search Engine Crawlers, SEO Crawlers, Uncategorized, and Undocumented AI Agents. This categorization allows us to analyze collective bot behaviors.
Ethical considerations. Data collection and analysis adhered to a study protocol approved by our local Institutional Review Board (IRB) [^4] IRB information redacted for anonymous submission. We followed all IRB guidelines for anonymizing and storing this dataset, including making significant efforts to exclude non-bot behavior from the dataset and preserve user privacy. We removed sessions originating from logged-in users from our institution or from institution-hosted networks, and anonymized IP addresses by hashing them. All data was stored and analyzed on secure servers only accessible by the research team.
@@ 3.2. Dataset overview
Most active bots. Now, we focus specifically on known bots, e.g. bots with self-declared user agents that are well-documented by several web sources. Table 3 gives details about the behavior of the 20 most active bots (as measured by number of web accesses) seen in our dataset. Together, these $20$ bots make up over $40\%$ of the observed web traffic. YisouSpider and AppleBot dominate, collectively driving $30\%$ of all traffic and scraping over $8$ GB of data during the 40 day period. YisouSpider is associated with the Chinese search engine Yisou and appears to be used to index web pages (Udger, 2025), specifically focusing on educational and job related webpages  (dar, 2025). We found that the vast majority of the YisouSpider’s accesses were to our institution’s people directory. AppleBot is a generic bot associated with Apple and is “used by Apple to index search results that allow the Siri AI Assistant to answer user questions” (Apple, 2025).
@@ 4.1. Methodology
Over a period of 8 weeks, we deployed four versions of robots.txt files on one of the $36$ institution-controlled sites we were monitoring. Data from this site was also part of the dataset in §3. This site was chosen because of its observed high bot traffic, making it a rich source of data on scraper behavior. Each robots.txt file was deployed for two weeks, then swapped out by support staff for the next version. The four versions of the robots.txt files have increasingly strict directives, ranging from no meaningful restrictions to full denial of access for certain bots. The restrictions were chosen to represent a steady gradient of robots.txt behaviors, ranging from permissive to strict while only changing one condition at a time, allowing us to assess how bot behaviors change in response to incremental robots.txt updates. We believe such controlled experimentation is the best approach, ensuring that confounding variables are not introduced by the modification of too many robots.txt properties at once. Using these related but distinct versions of robots.txt allowed us to achieve two distinct goals: measuring whether certain robots.txt directives were more or less respected by scrapers, and measuring how quickly scrapers adapted to new robots.txt restrictions. The files are described below and shown in Figures 8-8.
@@
  • Version 1 (Fig. 8): All bots retain the same level of access as in the base version, but a 30 second crawl delay is requested between successive page requests.
  • Version 2 (Fig. 8): Most bots are allowed to access only the /page-data/* endpoint, which we observed experimentally to be a common target for scrapers, but are disallowed from all other sites. Eight search engine optimization (SEO) bots [^5] Googlebot, Slurp, bingbot, Yandexbot, DuckDuckBot, BaiduSpider, DuckAssistBot, ia_archiver are exempt from these restrictions per our institution’s request, to ensure the sites remain easily findable online.
@@
Data preparation. We performed basic preprocessing on our datasets after collection to aid analysis. Much of our preprocessing pipeline for the robots.txt experiments mirrored that described in §3. However, for our analysis of compliance with robots.txt we also filter out bots that accessed the site less than $5$ times under any robots.txt version. Finally, we also eliminated any bots that appeared to have spoofed their user-agent. We determine potential spoofing by looking at bots for which $90\%$ of traffic originates from a single ASN, and flag as possibly spoofed instances any requests originating from another ASN (e.g. not the $90\%$ majority ASN). A deeper exploration of this spoofing metric can be found in §5.
@@
AI Assistants 0.910 (2077) 0.131 (2154) 1.000 (1126) 0.616
AI Data Scrapers 0.560 (1857) 0.352 (1824) 0.766 (1818) 0.559
AI Search Crawlers 0.895 (1157) 0.623 (974) 0.348 (1056) 0.631
Fetchers 0.925 (249) 0.283 (250) 0.377 (235) 0.531
Headless Browsers 0.036 (2777) 0.278 (4207) 0.011 (1282) 0.155
Intelligence Gatherers 0.809 (1120) 0.372 (549) 0.094 (1008) 0.450
Other 0.486 (5685) 0.139 (3801) 0.019 (3688) 0.255
SEO Crawlers 0.635 (1564) 0.831 (1215) 0.639 (1268) 0.695
Search Engine Crawlers 0.780 (6345) 0.366 (5925) 0.189 (6314) 0.447
Directive average 0.609 0.310 0.307 0.420
@@
Endpoint access compliance (v2). This robots.txt version restricts bots (except for a group of SEO bots, see §4.1) to only accessing subpages under a single endpoint, /page-data, on the site. Computing the ratio of compliance with this directive is more straightforward. For each unique user-agent, we count its accesses to either the robots.txt file (which is always allowed) or /page-data subpages and compute the ratio of this to the user agent’s total page accesses. This compliance ratio $C_{\tau_{i}}$ is $1$ if the user-agent only accesses the allowed pages, robots.txt and page-data, and is near 0 if it consistently accesses many other pages. We then compute an identical compliance metric for user agents under our default robots.txt and use this as the baseline for a paired z-test for differences in proportions between the two datasets. Ideally, we should see a significant uptick in compliance under the endpoint directive.